Math Geometry Question!!!!?
Evaluate this polynomial when x = -4 and y = 5
x^2y^2 - x^2 + y^2 - 3x + 2

Answer by algebrahouse.com(1659)

x²y² - x² + y² - 3x + 2, when x = -4 and y = 5
= (-4)²(5)² - (-4)² + (5)² - 3(-4) + 2 {substituted in}
= 16(25) - 16 + 25 + 12 + 2 {evaluated exponents and multiplied -3 by -4}
= 400 - 16 + 25 + 12 + 2 {multiplied 16 by 25}
= 423 {added and subtracted}

For more help from me, visit: www.algebrahouse.com